OpenStudy (anonymous):

For the following system of equations, what is the x-value of the solution? -x+2y=6 6y=x+18 A. -1 b.0 c.1 d.6

jhonyy9 (jhonyy9):

from first equation x=2y-6 - push this in second 6y=2y-6+18 4y=12 y=12/4 =3 x=2*3-6=6-6=0 x=0 y=3

OpenStudy (anonymous):

thankyou :)

jhonyy9 (jhonyy9):

yw. good luck bye